# The Quantum Decoding Problem : Tight Achievability Bounds and Application to Regev’s Reduction

### v1

@@ 4.2 A slight tweak in the algorithm which avoids measuring the zero codeword
Let $q\geqslant 2$ be a prime power. We assume that there exist three increasing sequences $(a_{i})_{i\geq 1}$, $(b_{i})_{i\geq 1}$, $(t_{i})_{i\geq 1}$ such that $a_{i}\geq b_{i}$ for every positive integer $i$ and the sequence $(1+t_{i}/a_{i})(1-t_{i}/b_{i})$ tends to a constant $H\in(0,1)$. Let $R$ be a constant in $(0,H)$. Consider a family $({\mathscr{C}}_{i})_{i\geq 1}$ of random linear codes where ${\mathscr{C}}_{i}$ is chosen with a generator matrix chosen uniformly at random in $\mathbb{F}_{q}^{\lfloor Rn_{i}\rfloor\times n_{i}}$. Then, for any $\epsilon>0$ the algorithm of the quantum reduction with PGM using the measurement basis $\{|Z_{{\boldsymbol{y}}}\rangle\}_{{\boldsymbol{y}}\in{\mathscr{S}}}$ described above, outputs codewords of $\mathscr{C}_{i}^{\perp}$ in the typical set ${\mathcal{T}}_{\epsilon}^{(n_{i})}$} with probability $1-\operatorname{negl}(i)$ as $i$ tends to infinity.
@@
$\leq$ $(1-p)\sqrt{\delta(\epsilon,n_{i})}$ (33)
@@
Let $q\geq 2$ be a prime power, $R$ be a fixed real number in $(0,1)$. We choose a random code $\mathscr{C}$} in $\mathbb{F}_{q}^{n}$ by choosing a generator matrix in ${\mathbb{F}_{q}}^{\lfloor Rn\rfloor\times n}$ uniformly at random. Let $p$ a probability distribution on $\mathbb{F}_{q}^{n}$. Let $\epsilon$ be a fixed positive real number. The probability over $\mathbf{G}$ that there exists a non zero codeword $\boldsymbol{c}$} in $\mathscr{C}^{\perp}$} such that $p({\boldsymbol{c}})>q^{-n(R-\epsilon)}$ is a negligible function of $n$.
@@
This lemma together with Proposition 7 leads for instance in the memoryless case to the following theorem.
@@ 4.3 Distribution associated to a metric
Let $F(\lambda)=\sum\limits_{\alpha\in{\mathbb{F}_{q}}}q^{-\lambda|\alpha|_{{\mathbb{F}_{q}}}}$. Observe that $F(0)=\sum\limits_{\alpha\in{\mathbb{F}_{q}}}q^{0|\alpha|_{{\mathbb{F}_{q}}}}=q$ and $\lim\limits_{\lambda\rightarrow\infty}F(\lambda)=0$. Since $F$ is continuous and decreasing over $0,\infty[$, we deduce that there exists a unique $\lambda>0$ such that $F(\lambda)=1$, which finishes the proof.

### v2

@@ 4.2 A tweak in the algorithm avoiding to measure the zero codeword
Let $q\geqslant 2$ be a prime power. We assume that there exist three increasing sequences $(a_{i})_{i\geq 1}$, $(b_{i})_{i\geq 1}$, $(t_{i})_{i\geq 1}$ such that $a_{i}\geq b_{i}$ for every positive integer $i$ and the sequence $(1+t_{i}/a_{i})(1-t_{i}/b_{i})$ tends to a constant $H\in(0,1)$. Let $R$ be a constant in $(0,H)$. Consider a family $({\mathscr{C}}_{i})_{i\geq 1}$ of random linear codes where ${\mathscr{C}}_{i}$ is chosen with a generator matrix chosen uniformly at random in $\mathbb{F}_{q}^{\lfloor Rn_{i}\rfloor\times n_{i}}$. Then, for any $\epsilon>0$ the algorithm of the quantum reduction with PGM using the measurement basis $\{|Y^{\prime}_{{\boldsymbol{y}}}\rangle\}_{{\boldsymbol{y}}\in{\mathscr{S}}}$ described above, outputs codewords of $(\mathscr{C}_{i}^{\perp})^{*}$ in the typical set ${\mathcal{T}}_{\epsilon}^{(n_{i})}$} with probability $1-\operatorname{negl}(i)$ as $i$ tends to infinity.
@@
$\leq$ $\frac{\sqrt{\delta(\epsilon,n_{i})}}{1-p}$ (41)
@@
Let $q\geq 2$ be a prime power, $R$ be a fixed real number in $(0,1)$. We choose a random code $\mathscr{C}$} in $\mathbb{F}_{q}^{n}$ by choosing a generator matrix in ${\mathbb{F}_{q}}^{\lfloor Rn\rfloor\times n}$ uniformly at random. Let $p$ be a probability distribution on $\mathbb{F}_{q}^{n}$. Let $\epsilon$ be a fixed positive real number. The probability over $\mathbf{G}$ that there exists a non zero codeword $\boldsymbol{c}$} in $\mathscr{C}^{\perp}$} such that $p({\boldsymbol{c}})>q^{-n(R-\epsilon)}$ is a negligible function of $n$.
@@
This proposition together with Proposition 7 leads for instance in the memoryless case to the following theorem.
@@ 4.3 Distribution associated to a metric
Let $F(\lambda)=\sum\limits_{\alpha\in{\mathbb{F}_{q}}}q^{-\lambda|\alpha|_{{\mathbb{F}_{q}}}}$. Observe that $F(0)=\sum\limits_{\alpha\in{\mathbb{F}_{q}}}q^{-0|\alpha|_{{\mathbb{F}_{q}}}}=q$ and $\lim\limits_{\lambda\rightarrow\infty}F(\lambda)=0$. Since $F$ is continuous and decreasing over $0,\infty[$, we deduce that there exists a unique $\lambda>0$ such that $F(\lambda)=1$, which finishes the proof.
@@
4.4 Solving the decoding problem in the surjective regime
As explained in the introduction a simple variation of Regev’s procedure allows to solve the decoding problem in the surjective regime. PGM can be used as for finding low weight codewords as follows
Algorithm 4.4 : Solving the decoding problem in the surjective regime with PGM. Input: $\mathscr{C}\subset\mathbb{F}_{q}^{n}$, $\boldsymbol{v}\in\mathbb{F}_{q}^{n}$
Output: an element of $\boldsymbol{c}$} such that $p({\boldsymbol{c}}-\boldsymbol{v})$ is large.
Initialization: $\quad|\phi_{0}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\sum_{\boldsymbol{e}\in\mathbb{F}_{q}^{n}}\chi_{\boldsymbol{c}}(-\boldsymbol{v})f(\boldsymbol{e})|{\boldsymbol{c}}\rangle|\boldsymbol{e}\rangle|{\boldsymbol{0}}\rangle$ adding $\boldsymbol{c}$} to $\boldsymbol{e}$: $\mapsto$ $\quad|\phi_{1}\rangle$ $=$ $\quad\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\chi_{\boldsymbol{c}}(-\boldsymbol{v})|{\boldsymbol{c}}\rangle\sum_{\boldsymbol{e}\in\mathbb{F}_{q}^{n}}f(\boldsymbol{e})|{\boldsymbol{c}}+\boldsymbol{e}\rangle|{\boldsymbol{0}}\rangle$ $=$ $\quad\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\chi_{\boldsymbol{c}}(-\boldsymbol{v})|{\boldsymbol{c}}\rangle|\psi_{\boldsymbol{c}}\rangle|{\boldsymbol{0}}\rangle$ apply the QFT on $|\psi_{\boldsymbol{c}}\rangle$: $\mapsto$ $\quad|\phi_{2}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\chi_{\boldsymbol{c}}(-\boldsymbol{v})|{\boldsymbol{c}}\rangle|\widehat{\psi_{\boldsymbol{c}}}\rangle|{\boldsymbol{0}}\rangle$ apply PGM coherently: $\mapsto$ $\quad|\phi_{3}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\chi_{\boldsymbol{c}}(-\boldsymbol{v})|{\boldsymbol{c}}\rangle\sum_{{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|Y_{{\boldsymbol{c}}^{\prime}}\rangle|{\boldsymbol{c}}^{\prime}\rangle$ $\text{ with }|\widehat{\psi_{\boldsymbol{c}}}\rangle=\sum_{{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|Y_{{\boldsymbol{c}}^{\prime}}\rangle$ erase $\boldsymbol{c}$}: $\mapsto$ $\quad|\phi_{4}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\chi_{\boldsymbol{c}}(-\boldsymbol{v})\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|{\boldsymbol{c}}-{\boldsymbol{c}}^{\prime}\rangle|Y_{{\boldsymbol{c}}^{\prime}}\rangle|{\boldsymbol{c}}^{\prime}\rangle$ reverse PGM: $\mapsto$ $\quad|\phi_{5}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\chi_{\boldsymbol{c}}(-\boldsymbol{v})\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|{\boldsymbol{c}}-{\boldsymbol{c}}^{\prime}\rangle|Y_{{\boldsymbol{c}}^{\prime}}\rangle|{\boldsymbol{0}}\rangle$ $=$ $\quad\sqrt{\textrm{P}_{\textrm{PGM}}}|{\boldsymbol{0}}\rangle\left(\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\chi_{{\boldsymbol{c}}}(-\boldsymbol{v})|Y_{\boldsymbol{c}}\rangle\right)+\sum_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}\neq{\boldsymbol{c}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|{\boldsymbol{c}}-{\boldsymbol{c}}^{\prime}\rangle\chi_{{\boldsymbol{c}}}(-\boldsymbol{v})|Y_{{\boldsymbol{c}}^{\prime}}\rangle$ meas. 1st register, if $|{\boldsymbol{0}}\rangle$: $\mapsto$ $\quad|\phi_{6}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\chi_{{\boldsymbol{c}}}(-\boldsymbol{v})|Y_{{\boldsymbol{c}}}\rangle=\dfrac{1}{Z_{\boldsymbol{v}{\mathbf{G}}^{\intercal}}}\sum_{{\boldsymbol{y}}\in\boldsymbol{v}+{\mathscr{C}}^{\perp}}\hat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle\;\;\text{w.p. $\textrm{P}_{\textrm{PGM}}$}$ meas. the whole state: $\mapsto$ $\quad|{\boldsymbol{y}}\rangle\;\;\text{with ${\boldsymbol{y}}\in\boldsymbol{v}+{\mathscr{C}}^{\perp}$.}$
The penultimate step (that is the first measuring step) can be justified by the following computation:
[…]
$=$ $\dfrac{1}{q^{k}\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{s}}\in\mathbb{F}_{q}^{k}}|\tilde{W_{\boldsymbol{s}}}\rangle\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\chi_{\boldsymbol{c}}(\boldsymbol{u}_{\boldsymbol{s}}-\boldsymbol{v}).$
It is readily verified that $\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\chi_{\boldsymbol{c}}(\boldsymbol{u}_{\boldsymbol{s}}-\boldsymbol{v})=0$ for all ${\boldsymbol{s}}\in\mathbb{F}_{q}^{k}$ with the exception of ${{\boldsymbol{s}}}^{\intercal}=\mathbf{G}{\boldsymbol{v}}^{\intercal}$, where $\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\chi_{\boldsymbol{c}}(\boldsymbol{u}_{\boldsymbol{s}}-\boldsymbol{v})=q^{k}$. This implies that
$\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\chi_{{\boldsymbol{c}}}(-\boldsymbol{v})|Y_{{\boldsymbol{c}}}\rangle$ $=$ $\dfrac{1}{\sqrt{|{\mathscr{C}}|}}|\widetilde{W_{\boldsymbol{v}{\mathbf{G}}^{\intercal}}}\rangle$.
By definition of $|\widetilde{W_{\boldsymbol{v}{\mathbf{G}}^{\intercal}}}\rangle$ we have
$|\widetilde{W_{\boldsymbol{v}{\mathbf{G}}^{\intercal}}}\rangle=\frac{1}{Z_{\boldsymbol{v}{\mathbf{G}}^{\intercal}}}\sum_{{\boldsymbol{y}}\in\boldsymbol{v}+{\mathscr{C}}^{\perp}}\hat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle,$
and this justifies that we get $\dfrac{1}{Z_{\boldsymbol{v}{\mathbf{G}}^{\intercal}}}\sum_{{\boldsymbol{y}}\in\boldsymbol{v}+{\mathscr{C}}^{\perp}}\hat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle$ after measuring $\boldsymbol{0}$} in the first register.
We have treated in the previous subsections the case ${\boldsymbol{s}}={\boldsymbol{0}}$ where we want to avoid to measure the trivial zero solution. This is why we used the projective measurement with respect to the $|Y^{\prime}_{\boldsymbol{y}}\rangle\langle Y^{\prime}_{\boldsymbol{y}}|$’s where $\boldsymbol{y}$} ranges over $\mathscr{S}$}. It is readily seen that Algorithm 4.4 can be analyzed in the same way we analyzed the modified version of Algorithm 4.1 based on the projective measurement with respect to the $|Y^{\prime}_{\boldsymbol{y}}\rangle\langle Y^{\prime}_{\boldsymbol{y}}|$’s. We replace in all the analysis $({\mathscr{C}}^{\perp})^{*}$ with $\boldsymbol{v}+{\mathscr{C}}^{\perp}$ and obtain readily a version of Theorem 9, Proposition 7, Proposition 8 where Algorithm 4.4 outputs elements of $\boldsymbol{v}+\mathscr{C}_{i}^{\perp}$ in the typical set ${\mathcal{T}}_{\epsilon}^{(n_{i})}$} with probability $1-\operatorname{negl}(i)$ as $i$ tends to infinity. There are also corresponding non achievability results, for instance a modified version of Proposition 9 showing that the probability over $\mathbf{G}$ that there exists an element $\boldsymbol{y}$} in $\boldsymbol{v}+{\mathscr{C}^{\perp}}$ such that $p({\boldsymbol{y}})>q^{-n(R-\epsilon)}$ is a negligible function of $n$. Again, this shows that at the limit where PGM works we basically get elements of $\boldsymbol{v}+{\mathscr{C}^{\perp}}$ of maximal probability $p$ with Algorithm 4.4. This gives in essence an operational version of [Ren18, §5, cor. 2]. We namely prove with this approach See 4